Circular No. 5678

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bn IN ANONYMOUS GALAXY Christian Pollas, Observatoire de la Côte d’Azur, reports his discovery

of an apparent supernova (mag ∼ 19) on technical pan films taken with the 0.9-m Schmidt telescope on Dec. 13.83 and 14.84 UT (limiting mag 21-22). SN 1992bn is located at α = 1^{h}26^{m}20^{s}.50, δ = +8^{o}53′24′′.8 (equinox 1950.0), or 1′′.2 east and 6′′.8 north of a galaxy of mag ∼ 16.5 having end figures 20^{s}.58, 31′′.6. Nothing is visible either on a similar film taken in 1989 or on the Palomar Sky Survey prints. A nearby star of m_{pg} ∼ 16 has end figures 26^{s}.26, 48′′.0.

NOVA SCORPII 1992 R. H. McNaught, Anglo-Australian Observatory, reports that, using

the coordinates by Kilmartin and Gilmore on IAUC 5528, he confirms the identification on the SERC J survey by Della Valle and Smette (IAUC 5529) of the pre-nova as a star of mag ∼ 18. Examination of several films in B, J, R, and I show the star to be blue with no evidence of variability. Coordinates measured from a J survey plate are α = 17^{h}03^{m}42^{s}.68, δ = −43^{o}11′26′′.3 (equinox 1950.0; uncertainty 0′′.3 in each coordinate), in close agreement with those on IAUC 5528.

NOVA SAGITTARII 1992 No. 3 McNaught also reports that the following improved coordinates were

measured from three CCD frames taken by himself on Oct. 14 using the Siding Spring 1.0-m telescope (the nova was better centered with respect to field stars than were the frames of Oct. 13; cf. IAUC 5635): α = 18^{h}20^{m}39^{s}.40, δ = −23^{o}01′05′′.2 (equinox 1950.0, uncertainty 0′′.2 in α and 0′′.5 in δ). The pre-nova has been identified as a blue star of mag ∼ 20.5 on the SERC J survey. It is faintly visible on the SERC and ESO R surveys, but not present on I survey plates. These I plates show surrounding stars that appear of similar brightness to the pre-nova on the J plates. Coordinates measured from one J plate have end figures 39^{s}.38, 05′′.5 (uncertainty 0′′.3 in each coordinate, reduced using the same 22 Hubble guide stars as for the CCD position above). Nearby stars have end figures 39^{s}.22, 07′′.6, mag 18; 39^{s}.22, 11′′.5, mag 17.

1992 December 19 Daniel W. E. Green